A 10kg box full of 1943 original Turtle Man comics falls from loading platform to a spring loaded platform 3 m below. Assume that there is no friction and that the spring can be modeled using Hooke's Law. Is the speed of the block at a minimum the instant the block strikes the spring? provide a calculation. If a 100kg box of Turtle Man action figures were placed on the spring platform, how much energy would be stored in the spring?


1
Expert's answer
2021-04-19T07:54:18-0400

Answer

Speed at the instant when strike the spring

"v=\sqrt{2gh}\\v=\sqrt{2(9.8) (3) }=7.67m/s"


Kinetic energy of block will be stored as

Potential energy of spring

"PE_{spring}=\frac{1}{2}mv^2"


"PE_{spring}=\frac{1}{2}(100)(7.67)^2=383.4J"